OpenStudy (anonymous):

If A and B are mutally exclusive events, P(A)=.26 and P(B)=.45 find: (a) P(not A) (b) P(AUB) (c)P(A intersection not B) (d)P(not A intersection not B)

OpenStudy (anonymous):

(a) P(not A) = 1-P(A) .45 = 1-.26 .45+.26 =.71 (b) P(AUB) not sure about this one. (c)P(A intersection not B) not sure about this one either (d)P(not A intersection not B) P(not A intersection not B)=1-P(AUB) 1=1-.71 =.29

OpenStudy (lgbasallote):

why do you have .45 = 1 - .26 for P(not A)?

OpenStudy (anonymous):

for \(P(A\cup B)\) add the probabilities since they are disjoint

OpenStudy (lgbasallote):

just so you know.... not A doesn't automatically mean B

OpenStudy (anonymous):

Oh okay, so would it just be 1-.26=.74 since P(not A)=1-P(A)?

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

So in that case would (c) be1-P(not(not A U not B))-P(not A intersection B) = 1-.29+.45 =.26?

OpenStudy (calculusfunctions):

NO!

OpenStudy (calculusfunctions):

@JerJason, Two events A and B are said to be mutually exclusive if\[P(A cupB)=P(A) + P(B)\]

OpenStudy (calculusfunctions):

That should be\[P(A \cup B)=P(A)+P(B)\]

OpenStudy (calculusfunctions):

If the events are mutually exclusive then\[P(A ∩ B)=0\]

OpenStudy (calculusfunctions):

Now use this information to answer your questions.